OpenStudy (anonymous):

calc 3 question! need help quickly!!! Find the potential function for F(x,y,z)= <(-cx/((x^2)+(y^2)+(z^2))), (-cy/((x^2)+(y^2)+(z^2))), (-cz/((x^2)+(y^2)+(z^2)))> Prove!

OpenStudy (anonymous):

Except at (0,0,0), the potential function is \[ -\frac{1}{2} c \log \left(x^2+y^2+z^2\right) \] Just take its gradient and see that is equal to your field.